can someone solve this question for me please?


the lenghth of one leg of a right triangle is 8 ft the lenghth of the hypotenuse is 4 feet longer than the other leg

Answer by nerdybill(7384)

the lenghth of one leg of a right triangle is 8 ft the lenghth of the hypotenuse is 4 feet longer than the other leg
.
Apply Pythagorean theorem:
Let x = length of other leg
then
8^2 + x^2 = (x+4)^2
8^2 + x^2 = (x+4)(x+4)
64 + x^2 = x^2+8x+16
64 = 8x+16
48 = 8
6 = x
Other leg is 6 feet
hypotenuse = x+4= 6+4 = 10 feet